3. Find the general solution of the differential equation y + 4y = 2 +e2x cos 2x

Find the general solution of the dierential equation
Answer #1

2x Note if t(o) y - P(x) Cos2x Given, Find the general soluation of the diffrential equation y + 4y = 2 te and if f(m)=o is asince D d ds d (2) = 0 dhe so B (2) = D(D (2)) = (6) dhe (2) = 0 YpCE (2- - Oto ) + €22. I Cosaa B²+ 4D+8 2x => Yp (2) = 1 /Since y (X) = G. Con2x + la sinza J, (x) = 1 / 4 (-2 Sin 22 – Cos 23) and 622 20 e = m + ( 25in 2x+ 2924) 2o so, the general

answered by: ANURANJAN SARSAM